KyledaGreat:

7y−4=−3(2−2x) Find the equation of the line which passes through the point (−12,−9) and is parallel to the given line. Express your answer in slope-intercept form. Simplify your answer.

KyledaGreat:

1 attachment
KyledaGreat:

@dukedadon

dukedadon:

interesting..

CripQUEZZ:

distribute first

KyledaGreat:

how ?

CripQUEZZ:

distribute the -3 into the parentheses

CripQUEZZ:

so itll be like -3 x 2 and -3 x -2x

KyledaGreat:

ok so like (-3 x 2) + (-3 x -2x) , correct me if i'm wrong

CripQUEZZ:

yes sir

CripQUEZZ:

so then ur problem will be 7y - 4 = -6 + 6x

surjithayer:

\[7y-4=-3(2-2x)\] 7y-4=-6+6x 7y=6x-6+4 7y=6x-2 \[y=\frac{ 6 }{ 7 }x-\frac{ 2 }{ 7 }\] its slope=6/7 eq. of line through (-12,-9) and slope 6/7 is \[y-(-9)=\frac{ 6 }{ 7 }(x-(-12))\] \[y+9=\frac{ 6 }{ 7 }(x+12)\] or 7y+63=6x+72 7y=6x+72-63 7y=6x+9

surjithayer:

or \[y=\frac{ 6 }{ 7 }x+\frac{ 9 }{ 7 }\]

KyledaGreat:

that's right the last one , thank you

KyledaGreat:

@surjithayer wrote:
or \[y=\frac{ 6 }{ 7 }x+\frac{ 9 }{ 7 }\]